Find the centroid of the area bounded by the parabola y=x2 and the line 2x+3.


1
Expert's answer
2021-08-02T14:48:45-0400

Let

"f(x) = x^2\ ,g(x) = 2x+3"

Formulas for centroid of area:

"A = \int_a^b(g(x)-f(x))dx\\
\overline x = \cfrac{1}{A} \int_a^b x(g(x)-f(x))dx\\
\overline y = \cfrac{1}{A} \int_a^b \cfrac{1}{2}((g(x))^2 - (f(x))^2)dx"

Find the "a" and "b":

"x^2 = 2x+3\\
x^2 -2x-3 = 0\\
x_1 = -1;x_2 = 3"

"A = \int_{-1}^3(2x+3-x^2)dx = x^2+3x-\cfrac{x^3}{3}|_{-1}^3 =\\
=9+9-9-(1-3-\cfrac{1}{3})=11+\cfrac{1}{3} = \cfrac{34}{3}"

"\overline x = \cfrac{3}{34}\int_{-1}^3 x(2x+3-x^2)dx = \cfrac{3}{34}\int_{-1}^3(2x^2+3x-x^3)dx =\\
=\cfrac{3}{34}(\cfrac{2x^3}{3}+\cfrac{3x^2}{2}-\cfrac{x^4}{4}|_{-1}^3)=\cfrac{3}{34}(18+\cfrac{27}{2}-\cfrac{81}{4}-(-\cfrac{2}{3}+\cfrac{3}{2}-\cfrac{1}{4})) = \cfrac{16}{17}"

"\overline y= \cfrac{3}{34}\int_{-1}^3\cfrac{1}{2}((2x+3)^2-x^4)dx = \cfrac{3}{68}\int_{-1}^3(4x^2+12x+9-x^4)dx = \\
= \cfrac{3}{68}(\cfrac{4x^3}{3}+6x^2+9x-\cfrac{x^5}{5})|_{-1}^3 = \cfrac{3}{68}(36+54+27-\cfrac{243}{5}-(-\cfrac{4}{3}+6-9+\cfrac{1}{5})) = \\
=\cfrac{16}{5}"

So centroid point's coordinate is "(\cfrac{16}{17};\cfrac{16}{5})"